OpenStudy (anonymous):

Select the correct product. (6 - 2x)(2x 2 - 1 + 3x)

OpenStudy (gorv):

expand

OpenStudy (gorv):

6*(2x 2 - 1 + 3x)-2x*(2x 2 - 1 + 3x) =12x2-6+18x-4x3+2x-6x2 =-4x3+12x2-6x2+2x+18x-6

OpenStudy (anonymous):

\[(6 - 2x)(2x^2 - 1 + 3x) =\] \[12x^2-6+18x -4x^3 +2x-6x^2 =-4x^3+6x^2+20x-6\]

OpenStudy (mathstudent55):

To multiply two polynomials together, multiply every term of the first polynomial by every term of the second polynomial. Then combine like terms and simplify.